OpenStudy (anonymous):

3(x+2)+4(x-3)=8 With the steps plz

OpenStudy (anonymous):

1. (3*x+2*3)+(4*x-4*3)=8 2. (3x+6)+(4x-12)=8 3. 3x+6+4x-12=8 4. 7x-6=8 {adding "+6" to both sides} 5. 7x=6+8 6. 7x=14 {dividing by "7" from both sides} 7. x=2